OpenStudy (anonymous):

Which of the following is the point-slope equation of the line with a slope of 2/5 and a point of (-3/6)? Could someone explain this in depth for future problems?

OpenStudy (anonymous):

Just to clarify, the point is (-3,6) right ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Okay point slope form is like this y – y1 = m(x – x1) where m is the slope, x1 is the x coordinate and y1 is the y coordinate, here we have the x1=-3, and y1=6, and m=2/5 to find the equation we plug these in to get y-6=2/5(x+3) make sense?

OpenStudy (anonymous):

Yes. It makes sense.

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

So the thing is, do I get y by itself first, or do I distribute 5 into X and 3?

OpenStudy (anonymous):

That is the equation in point-slope form, but if you wanted to simplify it more you would distribute 2/5 to the x and 3 then add 6

OpenStudy (anonymous):

I have to simplify it into y=mx+b form.

OpenStudy (anonymous):

Wait so it's y=2/5X+9?

OpenStudy (anonymous):

Hmm, not quite when you distribute you have to multiply the 2/5 by x and 3 what is 2/5*x+2/5*3

OpenStudy (anonymous):

1.2

OpenStudy (anonymous):

So you have y-6=2/5x+1.2 now add 6

OpenStudy (anonymous):

So I divide the (x+3) by 2/5? Then I add six? Are you sure...?

OpenStudy (anonymous):

*Multiply the (x+3)

OpenStudy (anonymous):

yes im sure

OpenStudy (anonymous):

Do I add six to everything? INcluding the 2/5X?

OpenStudy (anonymous):

I got y=6.4X+7.2

OpenStudy (anonymous):

I got it right. Thank you so much.